40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13998     Entry time: Thu Jun 21 15:32:05 2018
Author: gautam 
Type: Update 
Category: Electronics 
Subject: EX AA filter range change 

[steve, gautam]

I took this opportunity of EX downtime to change the supply voltage for the AA unit (4-pin LEMO front panel) in 1X9 from +/-5V to +/-15V. Inside the AA board are INA134 and DRV135 ICs, which are rated to work at +/-18V. In the previous state, the inputs would saturate if driven with a 2.5Vpp sine wave from a DS345 func. gen. After the change, I was able to drive the full range of the DS345 (10Vpp), and there was no saturation seen. This AA chassis is only used for the OSEM signals and also some ALS signals. Shadow sensor levels and spectra are consistent before and after the change. The main motivation was to not saturate the Green PDH Reflection signal in the digital readout. The steps we took were:

  1. Confirm (by disconnecting the power cable at the back of the AA box) that the power supplied was indeed +/- 5 V.
  2. Remove DIN fuse blocks from DIN rail for the relevant blocks.
  3. Identify a +15 V, -15 V and GND spot to plug the wires in. 
  4. Effect the swap.
  5. Re-insert fuses, checked supply voltage at connector end of the cable was now +/- 15 V as expected.
  6. Re-connect power cable to AA box.
ELOG V3.1.3-